# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ III Periodically driven harmonic oscillator
Figure 2: Flow trajectories of Eq. (8) for $\omega_{1}=\Omega$, (a) $\omega_{0}/\Omega=0$, and (b) $\omega_{0}/\Omega=0.5$ in the $(B_{0},B_{1})$ plane, depicted for varying $A/\Omega$. The $n-$th freezing point, denoted $\text{FP}_{n}$, corresponds to distinct values of $A/\Omega$ in (a) and (b), while $\text{NFP}_{n}$ stands for near-$\text{FP}_{n}$. ${\mathcal{H}}_{\text{pre}}$ denotes the subspace of prethermal Hamiltonians (locus $B_{1}=0$), and ${\mathcal{H}}_{\text{sym}}$ denotes subspace of $\rm{U}(1)$ symmetric Hamiltonians (locus $B_{0}=0$). Dynamical freezing corresponds to the flow landing exactly on the intersection of ${\mathcal{H}}_{\text{pre}}$ and ${\mathcal{H}}_{\text{sym}}$, as shown by the curves associated with $\text{FP}_{1}$ (red), $\text{FP}_{2}$ (orange), and $\text{FP}_{3}$ (brown).
@@
were $\hat{a}^{\dagger}(\hat{a})$ represents a bosonic creation (annihilation) operator, and $A,~{}\Omega$ represent the drive amplitude and frequency, respectively. Rewriting the above in terms of the decomposition in Eq. (1),
@@
Figure 3: (a) $P(\lambda)$ defined in Eq. (11) as a function of $A/\Omega$ for $\lambda\rightarrow\lambda_{c}$. The sharp drop in $P(\lambda_{c})$ marks the freezing points, FP1, FP2, and FP3, in the considered range of $A/\Omega$. The results for $P(\lambda)$ are obtained for system size $L=10$ using exact diagonalization, and for $L=40$ from MPO calculations. (b) $P(\lambda)$ as a function of fRG time, ($\lambda$), for the different freezing points shows a saturation to the plateau at “late” times. Results for all freezing points are shown for $L=40$ from MPO computations, and for non freezing point (NFP) for $L=10$ using ED. The drive frequency is fixed as $\Omega=10$. We show the results for $P(\lambda)$ for $L=10$ until $\lambda=10$ in the Appendix F
In this example, dynamical freezing is associated with an emergent ${\rm U}(1)$ symmetry generated by ${\cal{Q}}=\hat{a}^{\dagger}\hat{a}$, with the distance between $H_{0}(\lambda)$ and ${\mathcal{H}}_{\text{sym}}$ — the space of U(1)-symmetric Hamiltonians — measured by $|B_{0}(\lambda)|$. The flow in Fig. 2 follows a spiral trajectory in the $(B_{0},B_{1})$ plane. Since at the end of the flow, the oscillatory part of the Hamiltonian is completely removed, the flow will end on the $B_{1}=0$ line, the space of prethermal Hamiltonians ${\mathcal{H}}_{\text{pre}}$. For a non-freezing point, the flow ends on ${\mathcal{H}}_{\text{sym}}$ with $B_{0}(\lambda\to\infty)\neq 0$. This corresponds to the presence of a $\rm{U}(1)$ breaking term in the renormalized Hamiltonian. In contrast, freezing occurs for $B_{0}(\lambda\to\infty)=0$ (equivalently, $[H_{0}(\lambda\to\infty),{\cal{Q}}]=0$), linked to certain initial values of $A/\Omega$, which generally depend on $\omega_{0}$. In the limit $\omega_{0}/\Omega\to 0$, these values align with the Magnus expansion, namely, the zeros of the Bessel function $J_{0}(...)$ (see Sec. V and Appendix. C for more details). Since $H_{0}(\lambda\to\infty)$ depends on $A/\Omega$ continuously, the order of the freezing point can be related to the geometry of the flow trajectory. In particular, every time $H_{0}(\lambda)$ crosses the axis $B_{0}=0$, the order of the freezing point increases by one; see Fig. 2.
As a caveat, some of the features of the simple example shown above do not generalize to a generic many-body Hamiltonian. First, here the system is fully integrable and does not heat up in the conventional sense, meaning that the flow only explores the ramp-up and prethermal regimes, respectively. Second, since there is only one set of terms ($\hat{a},~{}\hat{a}^{\dagger}$) that violates the $\rm{U(1)}$ symmetry, we can realize freezing for any oscillator frequency $\omega_{0}/\Omega$ by tuning the parameter $A/\Omega$. In a generic many-body system, there can be multiple symmetry violating operators in $H_{0}(\lambda)$, and $[H_{0}(\lambda),{\mathcal{Q}}]$ need not vanish by tuning a single parameter $A/\Omega$. However, as we will show later in Sec. IV, the minimum value decays as a power-law in $1/\Omega$.
Figure 4: $P(\lambda\to\lambda_{c})$ for ($L=10,~{}20,~{}40$) at the first freezing point (FP1) exhibits a universal $1/\Omega^{2}$ scaling. Inset: $Q(\lambda=\lambda_{c})$ for different system sizes exhibits an exponential decay with increasing driving frequency. The data near FP2 also follows the same frequency scaling.
@@ IV Floquet spin chain: numerical results
We now turn to the main example of interest in this paper, and demonstrate that the key features discussed above can be generalized to a many-body setting using the same fRG. We have analyzed the flow trajectories for a one-dimensional, non-integrable Ising spin chain with a Floquet cosine-wave drive coupled to the transverse field. The results are sufficiently universal, even in the presence of additional perturbations to the Hamiltonian (see Appendix F for an additional example), that we believe our results are representative of the general phenomenology introduced above. The Hamiltonian is given by,
@@
where the second-neighbor coupling $J_{2}$ is responsible for the non-integrable nature; for $J_{2}=0$, freezing in the integrable (free) Hamiltonian has been analyzed previously [39]. We can split the above Hamiltonian into two parts,
$H_{0}(0)$ $=$ $\sum_{i}\left[-JS^{z}_{i}S^{z}_{i+1}-J_{2}S^{z}_{i}S^{z}_{i+2}\right]\,,$ (10a)
@@
which represent the initial condition for the flow associated with $H_{0}(\lambda)$ and $H_{\pm 1}(\lambda)$, respectively. Our numerical results will be presented in units where $J=1$. In Appendix F, we present results for $H_{0}(0)$ that includes an additional transverse magnetic field along the $x$ direction.
@@
which quantifies the extent to which the oscillatory component has been removed during the flow. As discussed in the discussion surrounding Eq. (5), a smaller $Q(\lambda)$ implies that $H_{0}(\lambda)$ is a better approximation for the effective Floquet Hamiltonian.
Figure 5: (a) Half chain entanglement entropy (EE) density, [i.e. $s_{1/2}=S_{1/2}/(L/2))$] plotted stroboscopically as a function of real time starting from a fully polarized state |$\ket{\psi}=\ket{x;+}^{\otimes L}$. The results are show for $L$=10 (using both exact and flow ED), and $L=40$ (using Flow MPO/MPS), at and away from freezing. Inset: Early time evolution of $s_{1/2}$ at FP1. Other parameters are set to $J=1$, $J_{2}=0.2$, and $\Omega=10$. (b) Frobenius norm of $H_{0}(\lambda)-H^{0}_{\text{magnus}}$ as a function of fRG time for varying $\Omega$. $H^{0}_{\text{magnus}}$ is computed analytically in Appendix E. Inset: Saturation value at late fRG time falls off as $1/\Omega^{2}$ at FP1 ($L=10$).
Figure 6: The Frobenius-norm of (a, b) $H_{1}$, and (c, d) $H_{0}$ as a function of fRG time for increasing $J_{2}$ for $L=10$. (a) Away from freezing, and for $J_{2}=0$, the system flows toward an integrable fixed point. For $J_{2}\neq 0$, at small fRG times the system first flows toward a prethermal fixed point, and eventually flows to a thermal fixed point. (b) At freezing, the system flows toward a prethermal fixed point for small and intermediate fRG times, controlled by the emergent approximate integrability, with the subsequent flow to the thermal fixed point delayed relative to (a). (c) Away from freezing, the onset of thermalization is controlled by the deviation away from the plateau in $H_{0}(\lambda)$, whose extent decreases with increasing $J_{2}$ (d) At freezing, the extent of the prethermal plateau in fRG time is enhanced compared to (c). Other parameters are set to: $\Omega=2J$, $J=1$, $A/\Omega=0.3$ away from freezing, $A/\Omega=0.601$ at FP1.
@@ IV.1 Approach to prethermal regime
To diagnose dynamical freezing numerically, we will evolve Eqs. (3) (i.e. let them flow) to a fRG time $\lambda=\lambda_{c}$, such that $Q(\lambda_{c})\ll 1$ and $P(\lambda)$ exhibits a $\lambda-$independent plateau. In other words, $\lambda_{c}$ is then in the initial prethermal regime. In this subsection, we choose $\lambda_{c}\lesssim\lambda_{\text{min}}$, the fRG time where $H_{1}(\lambda)$ minimizes. For all of the numerical computations in this section, we find that a $\lambda_{c}=1\times J^{-1}$ satisfies these numerical criteria. In general, we find that $\lambda_{c}$ should be set to the inverse intrinsic single-particle energy scales associated with the local thermalization dynamics.
@@
In Fig. 3(a), we show explicit results for $P(\lambda)$ at and away from the freezing points; clearly, this requires us to identify the location for the freezing points. Thus, we first performed a scan of $P(\lambda=\lambda_{c})$ over a fine interval of $A/\Omega$ (including for a $L=40$ chain using MPO), which clearly reveals the sharp features associated with freezing. These are encoded in the striking drop of $P(\lambda_{c})$ at specific values of $A/\Omega$. Note that $\lambda_{c}$ is chosen such that $P(\lambda)$ is already in a plateau-like regime, as explained above, and shown in Fig. 3(b). Interestingly, there is an “asymptotic” plateau associated with $P(\lambda)$ as a function $\lambda$, both at and away from freezing (We show in Appendix F that the plateau feature at $\Omega=10J$ can extend to much longer fRG time). However, the flow towards the plateau at the first, second, and third freezing points, are accompanied by a rich structure, reminiscent of some aspects of the flow trajectories for the driven harmonic oscillator in Fig. 2, which we shall describe in more detail below.
@@
The locations of the freezing points are roughly consistent with the expectations based on Floquet-Magnus expansion, namely $J_{0}(A/4\Omega)=0$, where $J_{0}$ is the Bessel function (see Appendix E for derivation). As we demonstrate in Appendix F, the position of the freezing point is renormalized away from the leading-order Magnus result, with a stronger renormalization for smaller $\Omega$.
@@
The flow initially shows a ramp-up behavior where $H_{0}(\lambda)$ (and relatedly, $P(\lambda)$) varies rapidly with the fRG time, $\lambda$. This is followed by the prethermal regime where $H_{0}(\lambda)$ flows slowly with $\lambda$, and as a result, $P(\lambda)$ reaches a plateau. As noted before, the hallmark of “dynamical freezing” is the strong reduction in the value associated with the plateau in $P(\lambda)$. It is worth highlighting that this plateau value at freezing, while at least an order of magnitude smaller than the initial value $P(0)$, does not vanish exactly. This leads to only an approximate conservation of $\mathcal{O}$.
@@
We note two additional features associated with $P(\lambda)$ at freezing. First, they develop strong “dips” in the ramp-up regime (marked by arrows in Fig. 3b), and the number of such dips is given by $(n-1)$, where $n$ represents the order of freezing point. This is reminiscent of the driven harmonic oscillator example in Fig. 2, where with varying $A/\Omega$, the flow trajectory can cross the ${\mathcal{H}}_{\text{sym}}$ plane, associated with the subspace of symmetric Hamiltonians. In the many-body situation, the crossing is not exact. Nevertheless associated with each such near crossing, the distance (in the space of Hamiltonians) between the effective Hamiltonian and ${\mathcal{H}}_{\text{sym}}$, which is given by $P(\lambda)$ develops a minimum. This corresponds to the dips. Second, for higher drive amplitudes $A/\Omega$, the flow takes a longer fRG time to reach the plateau. This can be intuitively understood as follows: the flow effectively integrates away the oscillatory part, $H_{1}(\lambda)$, at a rate proportional to the drive frequency $\Omega$, such that a stronger drive spends a longer time in the ramp-up regime. This feature is also consistent with our analytical results in Sec. V.
@@
It is worth emphasizing that the finite value of $P(\lambda)$ at freezing is in sharp contrast to the harmonic oscillator example in Fig. 2 presented earlier, where $P(\lambda)$ vanishes exactly. The difference can be traced back to our concluding comments in Sec. III, where we pointed out that in a many-body system, a variety of operators that violate the emergent conservation can be generated along the flow, which cannot all be suppressed by fine-tuning a single parameter, $A/\Omega$. However, these additional symmetry violating operators can be universally suppressed in the high frequency $\Omega\to\infty$ limit. In Fig. 4, we plot $P(\lambda_{c})$ at two distinct freezing points with increasing $\Omega$. Interestingly, we find that $P(\lambda_{c})$ decays as $1/\Omega^{2}$ at large $\Omega$, which is consistent with our analytical results in Sec. V. Relatedly, $Q(\lambda)$ (Eq. (12)) falls off exponentially with the drive frequency, $\Omega$ (Fig. 4 inset). We also point out that $P(\lambda)$, which is normalized to be an intensive quantity, already appears to show convergence to the thermodynamic limit for $L=20$, given the lack of any appreciable difference from the $L=40$ results.
@@
The results for $s_{1/2}$ obtained using both sets of methods are plotted in Fig. 5(a), which show excellent numerical agreement at and away from freezing. This provides further evidence for our Floquet flow renormalization being able to capture the essential universal aspects of the problem, including the growth of entanglement entropy in real-time. We also note that $H_{0}(\lambda)$ in the prethermal regime accurately captures the real-time dynamics up to an exponentially long time. Finally, we directly compare the “distance” between $H_{0}(\lambda)$ and $H^{0}_{\text{magnus}}$, the leading order result of the Floquet-Magnus expansion (see Appendix E for detail), which is expected to capture the prethermal dynamics in the $\Omega\to\infty$ limit. Indeed, we see that this distance decreases along the flow trajectory, $\lambda$, and reaches a plateau in the prethermal regime; the plateau value decays as $1/\Omega^{2}$ in the high frequency limit, which is consistent with our analytical result in Sec. V. Furthermore, we demonstrate in Appendix B that $H_{0}(\lambda_{c})$ is approximately isospectral to the actual Floquet Hamiltonian $H_{F}$.
Figure 7: (a) Frobenius norm of $H_{1}(\lambda)$ as a function of fRG time for $L=10$. Inset: isolated peak at late fRG time along with theoretical fit [$(a/2){\rm sech}(a(\lambda-b))$]. Parameters: $J=1$, $J_{2}=0.2$, $\Omega=2.0$ at FP1. (b) Frobenius norm of $H_{0}(\lambda)$, with a step-like feature correlated with the peak in $H_{1}(\lambda)$ in (a). Additional results for $L=6$ are shown in Appendix F
IV.2 Thermalization via instantons

### v2

@@ III Periodically driven harmonic oscillator
Figure 2: Flow trajectories of a driven harmonic oscillator given by Eq. (8) for $\omega_{1}=\Omega$, (a) $\omega_{0}/\Omega=0$, and (b) $\omega_{0}/\Omega=0.5$ in the $(B_{0},B_{1})$ plane, depicted for varying $A/\Omega$. The $n$-th freezing point, denoted $\text{FP}_{n}$, corresponds to distinct values of $A/\Omega$ in (a) and (b), while $\text{NFP}_{n}$ stands for near-$\text{FP}_{n}$. ${\mathcal{H}}_{\text{pre}}$ denotes the subspace of prethermal Hamiltonians (locus $B_{1}=0$), and ${\mathcal{H}}_{\text{sym}}$ denotes subspace of $\rm{U}(1)$ symmetric Hamiltonians (locus $B_{0}=0$). Dynamical freezing corresponds to the flow landing exactly on the intersection of ${\mathcal{H}}_{\text{pre}}$ and ${\mathcal{H}}_{\text{sym}}$, as shown by the curves associated with $\text{FP}_{1}$ (red), $\text{FP}_{2}$ (orange), and $\text{FP}_{3}$ (brown).
@@
[…]
@@
Figure 3: (a) $P(\lambda)$ defined in Eq. (12) as a function of $A/\Omega$ for $\lambda\rightarrow\lambda_{c}$. The sharp drop in $P(\lambda_{c})$ marks the freezing points, FP1, FP2, and FP3, in the considered range of $A/\Omega$. The results for $P(\lambda)$ are obtained for system size $L=10$ using exact diagonalization, and for $L=40$ from MPO calculations. (b) $P(\lambda)$ as a function of fRG time, ($\lambda$), for the different freezing points shows a saturation to the plateau at “late” times. Results for all freezing points are shown for $L=40$ from MPO computations, and for non freezing point (NFP) for $L=10$ using ED. Other parameters: $J=1$, $J_{2}=0.2$, $\Omega=10$. We show the results for $P(\lambda)$ for $L=10$ until $\lambda=10$ in the Appendix C.
[…]
As a caveat, some of the features of the simple example shown above do not generalize to a generic many-body Hamiltonian. First, here the system is fully integrable and does not “thermalize” in the conventional sense, meaning that the flow only explores the ramp-up and prethermal regimes, respectively. Second, since there is only one set of terms ($\hat{a},~\hat{a}^{\dagger}$) that violates the $\rm{U(1)}$ symmetry, we can realize freezing for any oscillator frequency $\omega_{0}/\Omega$ by tuning the parameter $A/\Omega$. In a generic many-body system, there can be multiple symmetry violating operators in $H_{0}(\lambda)$, and $H_{0}(\lambda),{\mathcal{Q}}]$ need not vanish by tuning a single parameter $A/\Omega$. However, as we will show later in Sec. IV, the minimum value decays as a power-law in $1/\Omega$. Moreover, many of the qualitative features discovered here via the flow trajectories generalize to varying degrees in the more generic setting, as we will highlight below.
@@ IV Floquet spin chain: numerical results
We now turn to the main example of interest in this paper, and demonstrate that the key features discussed above can be generalized to a many-body setting using the same fRG. We have analyzed the flow trajectories for a one-dimensional, non-integrable Ising spin chain with a Floquet cosine-wave drive coupled to the transverse field. The results are sufficiently universal, even in the presence of additional perturbations to the Hamiltonian, that we believe our results are representative of the general phenomenology introduced above. The Hamiltonian is given by,
@@
where the second-neighbor coupling $J_{2}$ is responsible for the non-integrable nature; for $J_{2}=0$, freezing in the integrable (free) Hamiltonian has been analyzed previously [21]. Note that in the absence of the drive term, the above Hamiltonian is entirely classical. Hence we have also analyzed the flow RG in the vicinity of freezing points as a function of a static transverse magnetic field in the $x$-direction, $B_{x}$, for the modified Hamiltonian,
$H(t)\rightarrow H(t)+\sum_{i}B_{x}S_{i}^{x}.$ (10)
We can split the above Hamiltonian into two parts,
$H_{0}(0)$ $=$ $\sum_{i}\left[-JS^{z}_{i}S^{z}_{i+1}-J_{2}S^{z}_{i}S^{z}_{i+2}+B_{x}S_{i}^{x}\right]\,$, (11a)
@@
which represent the initial condition for the flow associated with $H_{0}(\lambda)$ and $H_{\pm 1}(\lambda)$, respectively. Our numerical results will be presented in units where $J=1$. Since the phenomenology remains qualitatively similar with the inclusion of $B_{x}$, we first begin by setting $B_{x}=0$, and then study the problem as a function of increasing (but static) $B_{x}$. Nevertheless, there are interesting quantitative modifications to the freezing phenomenology with the inclusion of $B_{x}\neq 0$, which we will discuss briefly in Sec. IV.2 below, and elaborate further in greater detail in Appendix C
@@
which quantifies the extent to which the oscillatory component has been removed during the flow. As discussed in the discussion surrounding Eq. (5), a smaller $Q(\lambda)$ implies that $H_{0}(\lambda)$ is a better approximation for the effective Floquet Hamiltonian upto later times.
[…]
@@ IV.1 Approach to prethermal regime
To diagnose dynamical freezing numerically, we will evolve Eqs. (3) (i.e. let them flow) to a fRG time $\lambda=\lambda_{c}$, such that $Q(\lambda_{c})\ll 1$ and $P(\lambda)$ exhibits a $\lambda$-independent plateau, as introduced in regime (b) in Sec. II; see also Table 1. In other words, $\lambda_{c}$ is then in the initial prethermal regime. In this subsection, we choose $\lambda_{c}\lesssim\lambda_{\text{min}}$, the fRG time where $H_{1}(\lambda)$ minimizes. For all of the numerical computations in this section, we find that a $\lambda_{c}=1\times J^{-1}$ satisfies these numerical criteria. In general, we find that $\lambda_{c}$ should be set to the inverse intrinsic single-particle energy scales associated with the local thermalization dynamics.
@@
In Fig. 3(a), we show explicit results for $P(\lambda)$ for $B_{x}=0$ at and away from the freezing points; clearly, this requires us to identify the location for the freezing points. Thus, we first performed a scan of $P(\lambda=\lambda_{c})$ over a fine interval of $A/\Omega$ (including for a $L=40$ chain using MPO), which clearly reveals the sharp features associated with freezing. These are encoded in the striking drop of $P(\lambda_{c})$ at specific values of $A/\Omega$. Note that $\lambda_{c}$ is chosen such that $P(\lambda)$ is already in a plateau-like regime, as explained above, and shown in Fig. 3(b). Interestingly, there is an “asymptotic” plateau associated with $P(\lambda)$ as a function $\lambda$, both at and away from freezing (We show in Appendix C that the plateau feature at $\Omega=10J$ can extend to much longer fRG time). However, the flow towards the plateau at the first, second, and third freezing points, are accompanied by a rich structure, reminiscent of some aspects of the flow trajectories for the driven harmonic oscillator in Fig. 2, which we shall describe in more detail below. We note here that the data in Fig. 3(a) are obtained for a fixed $\Omega$ and varying $A$. The numerical results remain identical, if instead $A$ is fixed and $\Omega$ is varied (see Appendix C). However, the only requirement remains that both $A$ and $\Omega$ are large compared to the single particle energy scales in $H(t)$, with $A/\Omega\sim O(1)$ for FPn with $n\sim O(1)$.
@@
The locations of the freezing points are roughly consistent with the expectations based on Floquet-Magnus expansion, namely $J_{0}(A/4\Omega)=0$, where $J_{0}$ is the Bessel function (see Appendix F for derivation). As we demonstrate in Appendix C, the position of the freezing point is renormalized away from the leading-order Magnus result, with a stronger renormalization for smaller $\Omega$.
@@
The flow initially shows a ramp-up behavior where $H_{0}(\lambda)$ (and relatedly, $P(\lambda)$) varies rapidly with the fRG time, $\lambda$. This is followed by the prethermal regime where $H_{0}(\lambda)$ flows slowly with $\lambda$, and as a result, $P(\lambda)$ reaches a plateau. As noted before, the hallmark of “dynamical freezing” is the strong reduction in the value associated with the plateau in $P(\lambda)$. It is worth highlighting that this plateau value at freezing, while at least an order of magnitude smaller than the initial value $P(0)$, does not vanish exactly. This leads to only an approximate conservation of $\mathcal{O}$. In Sec. IV.2 we will return to this point when $B_{x}\neq 0$.
@@
We note two additional features associated with $P(\lambda)$ at freezing. First, they develop strong “dips” in the ramp-up regime (marked by arrows in Fig. 3b), and the number of such dips is given by (n1)(n-1), where $n$ represents the order of freezing point. This is reminiscent of the driven harmonic oscillator example in Fig. 2, where with varying $A/\Omega$, the flow trajectory can cross the ${\mathcal{H}}_{\text{sym}}$ plane, associated with the subspace of symmetric Hamiltonians. In the many-body situation, the crossing is not exact. Nevertheless associated with each such near crossing, the distance (in the space of Hamiltonians) between the effective Hamiltonian and ${\mathcal{H}}_{\text{sym}}$, which is given by $P(\lambda)$ develops a minimum. This corresponds to the dips. Second, for higher drive amplitudes $A/\Omega$, the flow takes a longer fRG time to reach the plateau. This can be intuitively understood as follows: the flow effectively integrates away the oscillatory part, $H_{1}(\lambda)$, at a rate proportional to the drive frequency $\Omega$, such that a stronger drive spends a longer time in the ramp-up regime. This feature is also consistent with our analytical results in Sec. VI.
@@
It is worth emphasizing that the finite value of $P(\lambda)$ at freezing is in sharp contrast to the harmonic oscillator example in Fig. 2 presented earlier, where $P(\lambda)$ vanishes exactly. The difference can be traced back to our concluding comments in Sec. III, where we pointed out that in a many-body system, a variety of operators that violate the emergent conservation can be generated along the flow, which cannot all be suppressed by fine-tuning a single parameter, $A/\Omega$. However, these additional symmetry violating operators can be universally suppressed in the high frequency $\Omega\to\infty$ limit. We find $P(\lambda_{c})$ scales with the frequency as $1/\Omega^{2}$ at the freezing points (see Appendix C for more detail).
@@
The results for $s_{1/2}$ obtained using both sets of methods are plotted in Fig. 4(a), which show excellent numerical agreement at and away from freezing. This provides further evidence for our Floquet flow renormalization being able to capture the essential universal aspects of the problem, including the growth of entanglement entropy in real-time. We also note that $H_{0}(\lambda)$ in the prethermal regime accurately captures the real-time dynamics up to an exponentially long time. It is worth noting that the growth of entanglement entropy in Fig. 4(a) shows clear differences between freezing points (FP1) and non-freezing points (NFP). Clearly at early to intermediate times, the freezing phenomenology is associated with much slower dynamics; however, this does not imply that this difference would survive up to much later times. The growth of entanglement is exponentially slow in the drive frequency as well, which can be seen quite clearly by focusing on smaller system sizes [1].
Finally, we directly compare the “distance” between $H_{0}(\lambda)$ and $H^{0}_{\text{magnus}}$ (quantified by Frobenius norm) in Fig. 4(b), the leading order result of the Floquet-Magnus expansion (see Appendix F for detail), which is expected to capture the prethermal dynamics in the $\Omega\to\infty$ limit. Indeed, we see that this distance decreases along the flow trajectory, $\lambda$, and reaches a plateau in the prethermal regime; the plateau value decays as $1/\Omega^{2}$ in the high frequency limit, which is consistent with our analytical result in Sec. VI. Furthermore, $H_{0}(\lambda_{c})$ is approximately isospectral to the actual Floquet Hamiltonian $H_{F}$ [1].
IV.2 Effect of a static transverse magnetic field
As noted previously, $H_{0}$ is purely classical when $B_{x}=0$; see Eq. (11a). Here, we analyze the effect of introducing an additional static transverse field along the $x$ direction. Specifically, we study the norm of $P(\lambda)$ as in Eq. (12) near the dynamical freezing point. In Fig. 5(a), we present results for $P(\lambda\to\lambda_{c})$ as a function of $(A/\Omega)$ near FP1 for increasing values of $B_{x}$. These results demonstrate that the location of the freezing point associated with the first (prethermal) fixed point is unaffected by increasing $B_{x}$, as indicated by the nearly identical locations of the minimum in $P(\lambda_{c})$ for all $B_{x}$.
[…]
$\textnormal{IPR}=\sum_{i_{x}=1}^{D_{H}}|\langle i_{x}|m\rangle|^{4},$ (14)
[…]
Figure 5: (a) $P(\lambda\to\lambda_{c})$ obtained from Eq. (12) as a function of $(A/\Omega)$ near FP1 with increasing $B_{x}$ for $L=10$. All other parameters are same as in Fig. 3. (b) IPR defined via Eq. (14) for eigenstates of $H_{0}(\lambda_{c})$ in the $\hat{S}_{x}$ basis plotted in decreasing order for $L=10$, comparing the freezing point (FP1) and the nonfreezing point (NFP), both with and without the magnetic field term in the $x$ direction. Inset: IPR vs $i/D_{H}$ for increasing system size for FP1 with $B_{x}\neq 0$. Other parameters are: $J=1$, $J_{2}=0.5$, $\Omega=5$, $\lambda_{c}=5.0$.
IV.3 Thermalization via instantons